# Can LLM Generate Culturally Relevant Commonsense QA Data? Case Study in Indonesian and Sundanese

### v2

@@ 6.4 Performance by Question Category
Figure 8 shows the LLMs’ performance across different question categories. It shows that LLMs perform better in the activity and place questions, which benefit from a lot of information readily available on the internet. However, they struggle with culinary questions, where specialized or cultural knowledge is often required. Interestingly, a significant improvement is seen from GPT-4 (0613 ver.) to GPT-4 Turbo (1106 ver.), particularly in Sundanese Human_Gen. This suggests that the model is acquiring more knowledge, which possibly comes from the interactions on the ChatGPT web. However, some categories still score below 80%, showing there is still room for improvement.
@@ 7 Discussion
Among the wrong answers, most cases come from overly general answers not specific to Indonesia, especially for questions in culinary, culture, and activity categories. For instance, given unique activity questions such as: "What is the mandatory song to be sung during the ‘mengheningkan cipta’ (moment of silence) in the flag ceremony?" the model gives a wrong answer in open-ended generation settings even if it could identify the correct answer in the multiple-choice setup. More failure examples can be seen in Table 13.
@@ 7.2 LLM-generated Data Quality
Due to the high performance of LLMs in answering LLM-generated data, one might argue that the data is less valuable. Ideally, creating a larger human-generated dataset is the most straightforward method to create a high-quality dataset. However, this approach may not always be feasible in the context of underrepresented languages, particularly given the limited resources. Building a dataset manually can be costly and challenging (we spent more than \$3,000 in total to build Human_Gen), and leveraging LLMs could be a practical and cheaper alternative, especially if we want to scale up the data.
To address the limitation of LLM-generated data, we can apply additional processes to reduce the noise. For instance, a collaboration between humans and models can be conducted to fix and revise the synthetic questions Liu et al. (2022); Putri and Oh (2022). This noise can also be reduced automatically, especially in Indonesian LLM_Gen questions, where errors typically involve minor typos. It is also important to note that not all LLM-generated data have low quality. Our experiments demonstrate that LLM_Gen has significantly fewer errors compared to LLM_Adapt, particularly in Sundanese. We also compared the performance between the cleaned vs. raw version of this data (Appendix F) and found that the performance variance between the original and cleaned data of LLM_Gen is relatively minor.
@@ 8 Conclusion
In this study, we introduced new Indonesian and Sundanese CommonsenseQA datasets built using various methods, including automatic dataset adaptation and generation with LLMs, as well as manual dataset creation by human annotators from diverse regions. Our thorough analysis reveals that adaptation from the existing English data is less effective for Sundanese. In contrast, when directly generated in the target languages, GPT-4 Turbo can produce relevant knowledge in both languages, although the cultural depth has not yet matched the human-generated ones. Nevertheless, given the limited resources, combining LLMs and humans for dataset creation can be a practical solution, allowing for a more efficient dataset creation by reducing the over-reliance on a single data source.
@@
All human-generated datasets have been manually validated to ensure that harmful or offensive questions are not present in the dataset. We also excluded potentially harmful questions in the LLM-generated datasets through automatic filtering. Our work has been reviewed by the Institutional Review Board (IRB) in our institution. All recruited annotators were paid above the minimum wage. Our datasets will be publicly available under the Creative Commons Non-Commercial (CC BY-NC 4.0) license.
@@ A.1 Prompts Details
A.1.1 Automatic Data Adaptation
@@
A.1.2 Automatic Data Generation
@@
Table 7 shows the win rate accuracy of concept, question, and options generation of Indonesian LLM, Merak-v4, and best-performing proprietary LLM, GPT-4 Turbo. The results demonstrate that GPT-4 Turbo significantly outperforms Merak-v4, with win rates ranging from 72% to 96%. Our sample-level analysis indicates that a significant number of questions generated by Merak-v4 tend to have obvious answers or are formulated as yes/no questions, such as "Apakah jagung dapat dimakan?" (eng: Is corn edible?). Additionally, despite being trained on Indonesian texts, the model occasionally produces questions with US-centric knowledge. For instance, it generates "Siapa yang memimpin kampanye pemilihan presiden pertama di Amerika Serikat?" (eng: Who led the first presidential election campaign in the United States?).
@@ A.3 Additional Analysis of the LLM-generated Data: Common Mistakes in Concept Adaptation
Status Num (%) Concept Example
Orig (ind) Modified (sun)
Correct 122 (77.22%)
hujan abu vulkanik
(volcanic ashfall)
hujan lebu vulkanik
(volcanic ashfall)
menara (tower) munara (tower)
Wrong 36 (22.78%)
cicak
(house gecko)
kadal imah
(house lizard)
klinik gigi
(dental clinic)
klinik dental
(dental clinic)
Table 8: Samples of Indonesian (ind) to Sundanese (sun) concept adaptation.
@@
Appendix B Annotators Demographics
@@
Appendix C Human Annotation Guideline
@@
Categories and Concepts
Five selected categories are used in the data creation process. Each category has 150 concepts that need to be used and developed to become commonsense questions. The following is a list of the categories.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
@@
There are several rules in the data creation process. The following is the explanation.
@@
Appendix D Manual Data Generation Findings
@@
As we employ annotators from different regions, it enables us to explore variations in the data they generate. We examine this by calculating the number of answer conflicts that arise during the "answering question" phase of our data generation pipeline (§4.2). The results from Indonesian and Sundanese annotators are detailed in Figure 5 and 6, respectively. Our analysis revealed that, out of the Indonesian annotators, the one from Bali has the highest number of conflicting answers. However, the number of questions generated by the Bali annotator does not seem to have a lot of conflicts, suggesting that the annotator tends to generate easier questions, some of which have an obvious answer. This finding contrasts with the results from West Java annotators, who, despite generating questions that lead to a higher number of conflicts, do so mainly due to the creation of more challenging option distractors.
In the case of Sundanese annotators, the variation in answer conflicts across regions is not significant. This lack of variation can likely be attributed to the Sundanese language’s narrower geographic distribution which primarily spoken in West Java, unlike the Indonesian language, which serves as a lingua franca within the country. This finding points to the fact that differences in annotators’ knowledge levels may arise from various factors, one of which could be the regional culture to which they are most exposed.
Appendix E Zero-Shot Benchmark Prompts
@@
Appendix F Effect of Synthetic Dataset Cleaning
@@ A. Gunung Jaya Wijaya
Appendix G LLM vs. Human-Generated Data Examples
@@
Appendix H Examples from Multiple-Choice vs. ‘Free’ Generation Experiment

### v3

@@ 6.4 Performance by Question Category
Figure 5 shows the LLMs’ performance across different question categories. It shows that LLMs perform better in the activity and place questions, which benefit from a lot of information readily available on the internet. However, they struggle with culinary questions, where specialized or cultural knowledge is often required. Interestingly, a significant improvement is seen from GPT-4 (0613 ver.) to GPT-4 Turbo (1106 ver.), particularly in Sundanese Human_Gen. This suggests that the model is “acquiring” more knowledge, possibly due to the interactions on the ChatGPT web. However, some categories still score below 80%, showing there is still room for improvement.
@@ 7 Discussion
Among the wrong answers, most cases come from overly general answers not specific to Indonesia, especially for questions in culinary, culture, and activity categories. For instance, given unique activity questions such as: "What is the mandatory song to be sung during the ‘mengheningkan cipta’ (moment of silence) in the flag ceremony?" the model gives a wrong answer in open-ended generation settings even if it could identify the correct answer in the multiple-choice setup. More failure examples can be seen in Table 13 in the Appendix.
@@ 7.2 LLM-generated Data Quality
Due to the high performance of LLMs in answering LLM-generated data, one might argue that such data is less valuable. Ideally, creating a larger human-generated dataset is the most straightforward method to ensure data quality. However, this approach may not always be feasible in the context of underrepresented languages, particularly given the limited resources. Building a dataset manually can be costly and challenging (we spent more than \$3,000 in total to build Human_Gen), and leveraging LLMs could be a practical and cheaper alternative, especially for scaling up the data.
To address the limitation of LLM-generated data, we can apply additional processes to reduce the noise. For instance, a collaboration between humans and models can be conducted to fix and revise the synthetic questions Liu et al. (2022); Putri and Oh (2022). This noise can also be reduced automatically, especially in Indonesian LLM_Gen questions, where errors typically involve minor typos.
It is also important to note that the quality of LLM-generated data can vary. Our experiments show LLM_Gen has significantly fewer errors than LLM_Adapt, particularly in Sundanese. We also compared the performance between the cleaned vs. raw version of the data (Appendix G) and found that cleaning the data had a relatively minor impact on performance in LLM_Gen. Still, it remains crucial not to depend solely on LLMs, as they still have limitations, such as producing convincing yet incorrect outputs Pan et al. (2023).
@@ 8 Conclusion
In this study, we introduced new Indonesian and Sundanese CommonsenseQA datasets built using various methods: automatic dataset adaptation, direct generation with LLMs, and manual dataset creation by human annotators from diverse regions. Our thorough analysis reveals that adapting existing English data is less effective for Sundanese. In contrast, direct generation in the target languages by GPT-4 Turbo produces relevant data for both languages, although the cultural depth has not yet matched the human-generated ones. Nevertheless, given the limited resources, combining LLMs and humans for dataset creation can be a practical solution, allowing for more efficient dataset creation by reducing the over-reliance on a single data source.
@@
Usage of GPT-4 for LLM_Gen
LLM_Gen can initially be seen as a dataset to “only” close the gap between local LLMs and GPT-4 since we solely use GPT-4 as the data generator. At the time of the experiment (October–December 2023), using GPT-4 was our only sufficient option. Moving forward, with the rise of better LLMs that support Indonesian, such as Command-R+ [^9] https://docs.cohere.com/docs/command-r-plus or Claude 3.5 Sonnet [^10] https://www.anthropic.com/news/claude-3-5-sonnet, we can enrich the existing data by including the generation of multiple “strong” LLMs. This will make LLM_Gen a more comprehensive testing tool. Additionally, we would like to emphasize that even if LLM_Gen is seen “only” as a tool for measuring the gap between local models and GPT-4, we believe its value should not be underestimated. Considering the lack of strong-performing open-source LLMs in low-resource languages like Sundanese, achieving comparable performance to GPT-4 in these languages would be highly valuable for the community.
Dataset Usage
Human_Gen and LLM_Gen datasets should be used exclusively as test data. When utilizing the training data portion in the LLM_Adapt dataset, special caution is required, particularly for the Sundanese language, where numerous errors have been identified. Training models on this data have a risk of error propagation of incorrect information. Additionally, the test data portions in both LLM_Adapt and LLM_Gen should be used carefully, as they contain “silver” labels and may include some inaccuracies. We strongly advise using the cleaned version of the data available on our GitHub repository and accompanying the evaluation results of the LLM-generated data with those from the Human_Gen dataset. This is because high performance of the evaluated model, especially on the LLM-generated data, does not necessarily indicate the robustness of the model. One should be careful when making claims.
@@
All human-generated datasets have been manually validated to ensure that harmful or offensive questions are not present in the dataset. We also excluded potentially harmful questions in the LLM-generated datasets through automatic filtering. Our work has been reviewed by KAIST Institutional Review Board (IRB). All recruited annotators were paid above the minimum wage based on the standard in Republic of Korea. Our datasets are publicly available under the Creative Commons Non-Commercial (CC BY-NC 4.0) license.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2024-00509258, AI Guardians: Development of Robust, Controllable, and Unbiased Trustworthy AI Technology).
@@ A.1 Prompts Details
A.1.1 Concept Relevancy Classifier
To classify whether a concept is relevant to the Indonesian and Sundanese context, we prompt GPT-3.5 Turbo [^11] We used gpt-3.5-turbo API checkpoint, accessed in September 2023. using an ensemble of five prompts. We condition the prompts based on whether the question concept is a verb or not, determined using the ConceptNet API. We set the location context as ‘Indonesia’ for Indonesian data and ‘West Java’ for Sundanese data. The examples of concept relevancy classifier results can be seen in Table 11.
Below are the prompts if the question concept is a verb (e.g., skiing, studying):
Prompt Variation 1
[…]
[leftmargin=0pt, rightmargin=0pt]
Text: {QUESTION}
Concept: {QUESTION CONCEPT}
Can one ’{QUESTION CONCEPT}’ in {Indonesia/West Java}? Answer with yes or no’.
[…]
Prompt Variation 2
[…]
[leftmargin=0pt, rightmargin=0pt]
Text: {QUESTION}
Concept: {QUESTION CONCEPT}
Do people in {Indonesia/West Java} familiar with ’{QUESTION CONCEPT}’ concept? Answer with yes or no’.
[…]
Prompt Variation 3
[…]
[leftmargin=0pt, rightmargin=0pt]
Text: {QUESTION}
Concept: {QUESTION CONCEPT}
Is ’{QUESTION CONCEPT}’ concept exist in {Indonesia/West Java}? Answer with yes or no’.
[…]
Prompt Variation 4
[…]
[leftmargin=0pt, rightmargin=0pt]
Text: {QUESTION}
Concept: {QUESTION CONCEPT}
In {Indonesia/West Java}, can people ’{QUESTION CONCEPT}’? Answer with yes or no’.
[…]
Prompt Variation 5
[…]
[leftmargin=0pt, rightmargin=0pt]
Text: {QUESTION}
Concept: {QUESTION CONCEPT}
Can ’{QUESTION CONCEPT}’ be done in {Indonesia/West Java}? Answer with yes or no’.
[…]
Below are the prompts if the question concept is not a verb (e.g., snow, bald eagle):
Prompt Variation 1
[…]
[leftmargin=0pt, rightmargin=0pt]
Text: {QUESTION}
Concept: {QUESTION CONCEPT}
Can one find ’{QUESTION CONCEPT}’ in {Indonesia/West Java}? Answer with yes or no’.
[…]
Prompt Variation 2
[…]
[leftmargin=0pt, rightmargin=0pt]
Text: {QUESTION}
Concept: {QUESTION CONCEPT}
Do people in {Indonesia/West Java} familiar with ’{QUESTION CONCEPT}’ concept? Answer with yes or no’.
[…]
Prompt Variation 3
[…]
[leftmargin=0pt, rightmargin=0pt]
Text: {QUESTION}
Concept: {QUESTION CONCEPT}
Is ’{QUESTION CONCEPT}’ concept exist in {Indonesia/West Java}? Answer with yes or no’.
[…]
Prompt Variation 4
[…]
[leftmargin=0pt, rightmargin=0pt]
Text: {QUESTION}
Concept: {QUESTION CONCEPT}
In {Indonesia/West Java}, can people find ’{QUESTION CONCEPT}’? Answer with yes or no’.
[…]
Prompt Variation 5
[…]
[leftmargin=0pt, rightmargin=0pt]
Text: {QUESTION}
Concept: {QUESTION CONCEPT}
Can ’{QUESTION CONCEPT}’ be found in {Indonesia/West Java}? Answer with yes or no’.
[…]
A.1.2 Automatic Data Adaptation
@@
A.1.3 Automatic Data Generation
@@
Table 8 shows the win rate accuracy of concept, question, and options generation of Indonesian LLM, Merak-v4, and best-performing proprietary LLM, GPT-4 Turbo. The results demonstrate that GPT-4 Turbo significantly outperforms Merak-v4, with win rates ranging from 72% to 96%. Our sample-level analysis indicates that a significant number of questions generated by Merak-v4 tend to have obvious answers or are formulated as yes/no questions, such as "Apakah jagung dapat dimakan?" (eng: Is corn edible?). Additionally, despite being trained on Indonesian texts, Merak-v4 occasionally produces questions with US-centric knowledge. For instance, it generates "Siapa yang memimpin kampanye pemilihan presiden pertama di Amerika Serikat?" (eng: Who led the first presidential election campaign in the United States?).
@@ A.3 Additional Analysis of the LLM-generated Data: Common Mistakes in Concept Adaptation
Appendix B Details on Categories and Concepts
Categories
Our LLM_Gen and Human_Gen datasets include five selected categories, detailed as follows:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
Concepts Creation
To create the concepts, we had a human annotator manually list concepts in the selected categories, ensuring they were culturally relevant but general enough for both Indonesian and Sundanese. For example, in the culinary category, concepts might include ‘duck’ or ‘lime leaves,’ avoiding specific dish names. We reviewed the concepts and provided feedback for revisions to ensure quality. This process results in 150 unique concepts per category in Indonesian. The annotators then manually translated these concepts into Sundanese during the QA pairs creation phase (§4.2). We aimed for parallel concepts to ensure a fair comparison between the languages.
Appendix C Annotators Demographics
@@
Appendix D Human Annotation Guideline
@@
There are several rules in the data creation process, detailed as follows:
@@
Appendix E Manual Data Generation Findings
@@
As we employ annotators from different regions, we can analyze variations in the data they generate. We examine this by calculating the number of answer conflicts that arise during the "answering question" phase of our data generation pipeline (§4.2). The results from Indonesian and Sundanese annotators are detailed in Figure 6 and 7, respectively.
Our analysis revealed that, out of the Indonesian annotators, the one from Bali has the highest number of conflicting answers. However, the number of questions generated by the Bali annotator does not seem to have a lot of conflicts, suggesting that this annotator tends to generate easier questions, some of which have an obvious answer. This finding contrasts with the results from West Java annotators, who, despite generating questions that lead to a higher number of conflicts, do so mainly due to the creation of more challenging option distractors.
Interestingly, in the case of Sundanese annotators, the variation in answer conflicts across regions is not significant. This lack of variation can likely be attributed to the Sundanese language’s narrower geographic distribution which primarily spoken in West Java, unlike the Indonesian language, which serves as a lingua franca within the country. However, it is important to note that given the small number of annotators, these trends cannot be conclusively linked to the annotators’ regional characteristics. The observed differences may also be due to individual differences rather than regional ones.
Appendix F Zero-Shot Benchmark Prompts
@@
Appendix G Effect of Synthetic Dataset Cleaning
@@ A. Gunung Jaya Wijaya
Appendix H LLM vs. Human-Generated Data Examples
@@
Appendix I Examples from Multiple-Choice vs. ‘Free’ Generation Experiment